# From affine algebraic racks to Leibniz algebras

### v1

@@ 6.2. Jacobi identity
To finish proving that $\mathfrak{q}$ is a Leibniz algebra, it remains to verify that $\cdot$$\cdot$] satisfies the Jacobi identity. Similarly to how the result for Lie racks relies on the left-distributivity axiom 2 of racks, our proof relies on the co-left distributivity identity 2 of corack algebras.
6.2.1.
@@
$S\coloneq((X\otimes Y\otimes Z)\circ(\mu\otimes\mathord{\operatorname{id}}\otimes\operatorname{id})\circ(\mathord{\operatorname{id}}\otimes\tau\otimes\operatorname{id})\circ(\nabla\otimes\nabla)\circ\nabla)(f)$
equals the sum of the two sums
@@
$=\sum[X,Y](u)(Z\circ(\varepsilon\otimes\operatorname{id})\circ\nabla)(v)$
@@
6.2.2.
Now we can finally finish proving that $\mathfrak{q}$ is a Leibniz algebra.
@@
The convolution operation $\cdot$$\cdot$] satisfies the Jacobi identity.
@@
$=(X\otimes Y\otimes Z)((\operatorname{id}\otimes\nabla)\circ\nabla)(f)$
@@
Corollary 6.7.
Viewing the adjoint representation of $\mathfrak{q}$ as a map $\operatorname{ad}\colon\mathfrak{q}\to\mathfrak{gl}(\mathfrak{q})$ as in Proposition 6.2, the image of $\operatorname{ad}$ lies in $\operatorname{Der}_{k}(\mathfrak{q},\mathfrak{q})$. That is, $X,\cdot]\in\operatorname{Der}_{k}(\mathfrak{q},\mathfrak{q})$ for all $k$-derivations $X\in\mathfrak{q}$.
@@ 6.3. Action on morphisms
Recall that if $(\mathfrak{q},[\cdot,\cdot]_{\mathfrak{q}})$ and $(\mathfrak{r},[\cdot,\cdot]_{\mathfrak{r}})$ are Leibniz algebras over $k$, then a $k$-linear map $\psi\colon\mathfrak{q}\to\mathfrak{r}$ is called a Leibniz algebra homomorphism if it preserves the respective Leibniz brackets. That is,
$\psi([X,Y]_{\mathfrak{q}})=[\psi(X),\psi(Y)]_{\mathfrak{r}}$
for all vectors $X,Y\in\mathfrak{q}$.
@@
The assignments $Q\mapsto\mathfrak{q}$ and $\varphi\mapsto\varphi^{\prime}\coloneq(d\varphi)_{e}$ define a covariant functor ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})\to{\mathsf{Leib}_{k}}$.
@@
Functoriality follows from the chain rule. Therefore, we only have to show that if $\varphi\colon Q\to R$ is a morphism of affine algebraic racks, then $\varphi^{\prime}$ is a Leibniz algebra homomorphism $\mathfrak{q}\to\mathfrak{r}$, where $(\mathfrak{q},[\cdot,\cdot]_{\mathfrak{q}})$ and $(\mathfrak{r},[\cdot,\cdot]_{\mathfrak{r}})$ are the Leibniz algebras of $Q$ and $R$, respectively.
@@
for all $k$-derivations $D\in\mathfrak{q}$. Certainly, $\varphi^{\prime}$ is $k$-linear. That $\varphi^{\prime}(\mathfrak{q})\subseteq\mathfrak{r}$ follows from the identification of $\mathfrak{q}$ and $\mathfrak{r}$ as tangent spaces at the identity via Proposition 5.2 (though verifying this directly is straightforward). Thus, it only remains to verify that $\varphi^{\prime}$ preserves the Leibniz bracket. Recall that
@@
Isomorphic affine algebraic racks have isomorphic Leibniz algebras.
@@
7. Applications of Theorem 1.1
In this section, we discuss several examples and properties of Leibniz algebras of affine algebraic racks. We use Theorem 1.1 throughout.
7.1. Omni-linear rack schemes
To illustrate the construction in Theorem 1.1, we compute the Leibniz algebra of the omni-linear rack scheme $\operatorname{OL}_{n}$. This provides a class of affine algebraic racks whose Leibniz algebras are not Lie algebras.
7.1.1.
First, we recall a construction of Weinstein [weinstein] in 2000 in the form studied by Liu and Sheng [omni]. Given a vector space $V$, define a new vector space $\mathfrak{ol}(V)\coloneq\mathfrak{gl}(V)\oplus V$. (In particular, $\mathfrak{ol}(V)$ has the same underlying vector space as the Lie algebra $\mathfrak{aff}(V)$ of the group $\mathrm{Aff}(V)$ of affine transformations of $V$.) Define a Leibniz bracket on $\mathfrak{ol}(V)$ by
$(X,v),(Y,w)]\coloneq(XY-YX,Xw).$
Then $\mathfrak{ol}(V)$ is a hemisemidirect product Leibniz algebra (see [kinyon2]) called an omni-Lie algebra. Despite its name, $\mathfrak{ol}(V)$ is a Lie algebra if and only if $\dim(V)=0$. (See [weinstein] for the original motivation for this name; cf. [omni]*Prop. 3.1.) If $V=k^{n}$, then we denote $\mathfrak{ol}_{n}(k)\coloneq\mathfrak{ol}(k^{n})$.
7.1.2.
We analogize a calculation of Kinyon [kinyon] in the algebraic setting. Recall that, as an affine scheme, the omni-linear rack scheme is
$\operatorname{OL}_{n}=\operatorname{GL}_{n}\times\mathbb{A}^{n}.$
Therefore, as a vector space, the tangent space of $\operatorname{OL}_{n}$ at the identity $e=(I_{n},\mathbf{0})$ is
$T_{e}(\operatorname{OL}_{n})=\mathfrak{gl}_{n}(k)\oplus k^{n}=\mathfrak{ol}_{n}(k).$
Recall from (4.1) that $\mathcal{O}(\operatorname{OL}_{n})$ is generated by the subset $\{s_{11},s_{12},\dots,s_{nn},{\det}^{-1},t_{1},\dots,t_{n}\}$ as a $k$-algebra. As a $k$-derivation $\mathcal{O}(\operatorname{OL}_{n})\to k$, each element $(X,v)\in T_{e}(\operatorname{OL}_{n})$ acts on $\mathcal{O}(\operatorname{OL}_{n})$ by
Proposition 7.1.
For all $n\geq 0$, the Leibniz algebra of the omni-linear rack scheme $\operatorname{OL}_{n}$ is the omni-Lie algebra $\mathfrak{ol}_{n}(k)$.
Proof.
By Remark 6.3, the formula
$(A,v)\triangleright(B,w)=(ABA^{-1},Aw)$
To compute the value of the Leibniz bracket on the second direct summand $k^{n}\leq T_{e}(\operatorname{OL}_{n})$, apply the convolution formula in Theorem 1.1 and then use (4.2) to write
$[(X,v),(Y,w)](t_{k})$ $=((X,v)\otimes(Y,w))(\nabla t_{k})$
$=((X,v)\otimes(Y,w))\left(\sum^{n}_{i=1}s_{ki}\otimes t_{i}\right)$
$=\sum^{n}_{i=1}x_{ki}w_{i}=(Xw)_{k}$
for all $(X,v),(Y,w)\in T_{e}(\operatorname{OL}_{n})$ and $1\leq k\leq n$. Hence, the Leibniz bracket of $T_{e}(\operatorname{OL}_{n})$ is given by
[(X,v),(Y,w)]=(XYYX,Xw),[(X,v),(Y,w)]=(XY-YX,Xw),
which is precisely the Leibniz bracket of $\mathfrak{ol}_{n}(k)$. ∎
Remark 7.2.
Comment 7.3.
In 2024, Liu and Sheng [omni] introduced omni-representations of Leibniz algebras, which are homomorphisms into the omni-Lie algebra $\mathfrak{ol}(V)$. In particular, every Leibniz algebra $\mathfrak{q}$ embeds into $\mathfrak{ol}(\mathfrak{q})$ via the adjoint omni-representation $X\mapsto([X,-],X)$; see [kinyon2]*Cor. 3.2 or [omni]*Prop. 3.1 for further discussion. Since the underlying vector space of $\mathfrak{ol}(V)$ is isomorphic to the Lie algebra $\mathfrak{aff}(V)$ of the group of affine transformations $\mathrm{Aff}(V)$, it follows that every finite-dimensional Leibniz algebra can be realized as a matrix Leibniz algebra.
[…]
7.2. From rack structure to Leibniz algebra structure
In this section, we provide several results relating the structure of an affine algebraic rack to the structure of its Leibniz algebra. These results recover the analogous results for the Lie algebras of affine algebraic groups. In the following, let $Q=\operatorname{Spec}(A)$ and $R=\operatorname{Spec}(B)$ be affine algebraic racks with Leibniz algebras $\mathfrak{q}$ and $\mathfrak{r}$, respectively.
7.2.1.
The following two results relate closed subracks and left ideals of affine algebraic racks to Leibniz subalgebras and left ideals of their Leibniz algebras, respectively. These results are shown more or less identically to the analogous results for Lie algebras of affine algebraic groups, but we provide the proofs for the reader’s convenience.
Recall that if $(\mathfrak{g},[\cdot,\cdot])$ is a Leibniz algebra, then a subspace $\mathfrak{h}\leq\mathfrak{g}$ is called a Leibniz subalgebra if $\mathfrak{h},\mathfrak{h}]\subseteq\mathfrak{h}$ and a left ideal if $\mathfrak{g},\mathfrak{h}]\subseteq\mathfrak{h}$. On the other hand, we say that $\mathfrak{g}$ is abelian if $\mathfrak{g},\mathfrak{g}]=0$.
@@ 7.2. From rack structure to Leibniz algebra structure
7.2.2.
Recall that the left center of a Leibniz algebra $(\mathfrak{g},[\cdot,\cdot])$ is the subset
@@
which is an abelian left ideal of $\mathfrak{g}$. [^5] One may analogously define a right center $Z_{r}(\mathfrak{g})$, but $Z_{r}(\mathfrak{g})$ is not a Leibniz subalgebra of $\mathfrak{g}$ in general. Note that if $\mathfrak{g}$ is a Lie algebra, then $Z_{\ell}(\mathfrak{g})$ is just the center of $\mathfrak{g}$ as a Lie algebra.
The following is a consequence of Propositions 4.10, 6.2, and 7.5.
@@
Corollary 7.11 (and, hence, Proposition 7.9) fails if the assumption that $\operatorname{char}(k)\neq 2$ is dropped. For example, let $G$ be the Heisenberg group. If $\operatorname{char}(k)=2$, then $g^{2}\in Z(G)$ for all $g\in G$. However, the Lie algebra of $G$ is the $3$-dimensional Heisenberg Lie algebra, which is not abelian. The converse of Corollary 7.11 also does not hold; for example, see [milne]*10.35.
Finally, combining Propositions 4.1, 4.7, and 4.8 with Example 7.10 yields the following.

### v2

@@ 8.2. Jacobi identity
To finish the proof of Theorem 1.1, it remains to verify that $\cdot$$\cdot$] and $\{\cdot,\cdot\}$ satisfies the left and right Jacobi identities, respectively.
Similarly to how the analogous result for Lie racks [kinyon]*Thm. 3.4 relies on the left-distributivity axiom 2 of racks, our verification of the left Jacobi identity for $\cdot$$\cdot$] relies on the co-left distributivity identity 2 and the co-fixing identity 4. The verification of the right Jacobi identity for $\{\cdot,\cdot\}$ is similar; we leave it to the reader to adapt the argument using the co-right distributivity identity 3 and the co-inverse fixing identity 4.
8.2.1.
@@
$S\coloneq((X\otimes Y\otimes Z)\circ(\mu\otimes\mathord{\mathrm{id}}\otimes\mathrm{id})\circ(\mathord{\mathrm{id}}\otimes\tau\otimes\mathrm{id})\circ(\nabla\otimes\nabla)\circ\nabla)(f),$
which comes from 2, can be written as the sum of two expressions
@@
$=\sum[X,Y](u)(Z\circ(\varepsilon\otimes\mathrm{id})\circ\nabla)(v)$
@@
8.2.2.
Now we can finally complete the proof of Theorem 1.1.
@@
The convolution operation $\cdot$$\cdot$] satisfies the left Jacobi identity.
@@
$=(X\otimes Y\otimes Z)((\mathrm{id}\otimes\nabla)\circ\nabla)(f)$
@@ 8.3. Action on morphisms
The assignments $Q\mapsto(\mathfrak{q},[\cdot,\cdot])$ (resp. $Q\mapsto(\mathfrak{q},\{\cdot,\cdot]\}$) and $\varphi\mapsto\varphi^{\prime}\coloneq(d\varphi)_{e}$ define a covariant functor ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})\to{\mathsf{Leib}_{k}}$ (resp. ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})\to{\mathsf{RLeib}_{k}}$).
@@
Once again, we prove the claim for ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})\to{\mathsf{Leib}_{k}}$; the proof for ${\mathsf{Rack}_{\ast}}({\mathsf{Aff}_{k}})\to{\mathsf{RLeib}_{k}}$ is similar.
Functoriality follows from the chain rule. Therefore, we only have to show that if $\varphi\colon Q\to R$ is a morphism of affine algebraic racks, then $\varphi^{\prime}$ is a Leibniz algebra homomorphism $\mathfrak{q}\to\mathfrak{r}$, where $(\mathfrak{q},[\cdot,\cdot]_{\mathfrak{q}})$ and $(\mathfrak{r},[\cdot,\cdot]_{\mathfrak{r}})$ are the left Leibniz algebras of $Q$ and $R$, respectively.
@@
for all $k$-derivations $D\in\mathfrak{q}$. Certainly, $\varphi^{\prime}$ is $k$-linear. That $\varphi^{\prime}(\mathfrak{q})\subseteq\mathfrak{r}$ follows from the identification of $\mathfrak{q}$ and $\mathfrak{r}$ as tangent spaces at the identity via Proposition 7.3 (though verifying this containment directly is straightforward).
Thus, it only remains to verify that $\varphi^{\prime}$ preserves the Leibniz bracket. Recall that
@@
The left and right Leibniz algebras $(\mathfrak{q},[\cdot,\cdot]),(\mathfrak{q},\{\cdot,\cdot\})$ are invariants of affine algebraic racks.
@@
9. From rack structure to Leibniz algebra structure
As applications of Theorem 1.1, we provide several results relating the structure an affine algebraic rack to the structure of its left Leibniz algebra. These results recover the analogous results for the Lie algebras of affine algebraic groups. We leave it to the reader to adapt these results to the right Leibniz algebras of affine algebaic racks.
In the following, let $Q=\operatorname{Spec}(A)$ and $R=\operatorname{Spec}(B)$ be affine algebraic racks with left Leibniz algebras $\mathfrak{q}$ and $\mathfrak{r}$, respectively.
9.1. Subracks and ideals
The following results relate closed subracks and left ideals of affine algebraic racks to Leibniz subalgebras and left ideals of their left Leibniz algebras, respectively. These results are shown more or less identically to the analogous results for Lie algebras of affine algebraic groups, but we provide the proofs for the reader’s convenience.
9.1.1.
Recall that if $(\mathfrak{g},[\cdot,\cdot])$ is a left Leibniz algebra, then a subspace $\mathfrak{h}\leq\mathfrak{g}$ is called a Leibniz subalgebra if $\mathfrak{h},\mathfrak{h}]\subseteq\mathfrak{h}$ and a left ideal if $\mathfrak{g},\mathfrak{h}]\subseteq\mathfrak{h}$. On the other hand, we say that $\mathfrak{g}$ is abelian if $\mathfrak{g},\mathfrak{g}]=0$.
@@ 9.1. Subracks and ideals
9.1.2.
Recall that the left center of a left Leibniz algebra $(\mathfrak{g},[\cdot,\cdot])$ is the subset
@@
which is an abelian left ideal of $\mathfrak{g}$. [^5] One may analogously define a right center $Z_{r}(\mathfrak{g})$, but $Z_{r}(\mathfrak{g})$ is not a left Leibniz subalgebra of $\mathfrak{g}$ in general. Note that if $\mathfrak{g}$ is a Lie algebra, then $Z_{\ell}(\mathfrak{g})$ is just the center of $\mathfrak{g}$ as a Lie algebra.
The following is a consequence of Propositions 6.11, 8.2, and 9.2.
@@ 9.2. Abelian conditions
The second part of Corollary 9.8 (and, hence, Proposition 9.6) fails if the assumption that $\operatorname{char}(k)\neq 2$ is dropped. For example, let $G$ be the Heisenberg group. If $\operatorname{char}(k)=2$, then $g^{2}\in Z(G)$ for all $g\in G$. However, the Lie algebra of $G$ is the $3$-dimensional Heisenberg Lie algebra, which is not abelian. The converse of Corollary 9.8 also does not hold; for example, see [milne]*10.35.
Finally, combining Propositions 6.1, 6.8, and 6.9 with Example 9.7 yields the following.